quadratic function: f
whose zeros are
-1
and
6
.

Answer by Edwin McCravy(20055)


If you were asked to solve such a quadratic function,
your last step would have been

  x = -1;    x = 6

Then the step before that would have been:

x + 1 = 0;   x - 6 = 0

Then the step before that would have been:

(x + 1)(x - 6) = 0

Then the step before that would have been:

x² - 5x - 6 = 0

Then the step before that would have been:

f(x) = x² - 5x - 6

Edwin